The total cost function of a monopolistic producer of two goods is TC = 3x + xy + 4y, where x and y denote the number of units of good 1 and good 2, respectively.


If p1 and p2 denote the corresponding prices, then the demand function of each good is p1 = 60 − x + y and p2 = 40 + 2x − y. Find the maximum profit if the firm is contracted to produce a total of 200 goods.



1
Expert's answer
2021-12-22T13:55:54-0500

"TC = 3x + xy + 4y"

"P_1 = 60 \u2212 x + y" and "P_2 = 40 + 2x \u2212 y"

"\pi=xp_1+yp_2-(3x+xy+4y)"

"=x(60-x+y)+y(40+2x-y)-(3x+xy+4y)"

"=60x-x^2+xy+40y+2xy-y^2-3x-xy-4y"

"=-x^2+2xy+57x-y^2+36y"


FOC:

"D\pi(x,y)=\begin{bmatrix}
 \frac{\Delta\pi}{ \Delta x} \\
 \frac{\Delta\pi}{ \Delta y}
\end{bmatrix}"



"=\begin{bmatrix}
 -2x+2y+57 \\
 2x -2y+36
\end{bmatrix}"


Solving simultaneously;

"x=5.25\ \\y=2.13"


"P_1 = 60 \u2212 5.25 + 2.13=56.88" and "P_2 = 40 + 2(5.25)-2.13=48.37"


Maximum profit;

"\pi=xp_1+yp_2-(3x+xy+4y)"

"=5.25(56.88)+2.13(48.37)-3(5.25)+5.25\times2.13+4(2.13)=405.6"